How many grams of mercuric chloride must react to give 5.81 g of Hg?
Co(s)+HgCl2(aq)→CoCl3(aq)+Hg(l)
Express your answer with the appropriate units.
1
Expert's answer
2020-12-01T14:24:09-0500

Balanced equation : 2Co + 3HgCl2 → 2CoCl3 + 3Hg

3 moles of mercuric chloride = 3 moles of Hg

3 ( 271.6g) mercurial chloride = 3 (200.6)g Hg

1g Hg= 3 × 271.6 /(3×200.6)g mercuric chloride

5.81g Hg = 3 ×271.6 ×5.81/(3×200.6)g mercuric chloride

= 7.87g mercuric chloride